### v1

# Half-integer thermal conductance in the absence of Majorana mode

Abstract
Considering a range of candidate quantum phases of matter, half-integer thermal conductance ($\kappa_{\text{th}}$) is believed to be an unambiguous evidence of non-Abelian states. It has been long known that such half-integer values arise due to the presence of Majorana edge modes, representing a significant step towards topological quantum computing platforms. Here we break this long-standing paradigm, reporting a comprehensive theoretical and experimental study where half-integer two-terminal thermal conductance plateau is realized employing Abelian phases. Our proposed setup features a confined geometry of bilayer graphene, interfacing distinct particle-like and hole-like integer quantum Hall states. Each segment of the device exhibits full charge and thermal equilibration. Our approach is amenable to generalization to other quantum Hall platforms, and may give rise to other values of fractional (electrical and thermal) quantized transport. Our study demonstrates that the observation of robust non-integer values of thermal conductance can arise as a manifestation of mundane equilibration dynamics as opposed to underlying non-trivial topology.
{affiliations}
Department of Physics, Indian Institute of Science, Bangalore, 560012, India.
Department of Condensed Matter Physics, Weizmann Institute of Science, Rehovot 76100, Israel.
National Institute of Material Science, 1-1 Namiki, Tsukuba 305-0044, Japan.
Department of Physics, Indian Institute of Science Education and Research (IISER) Tirupati, Tirupati 517619, India.
Raymond and Beverly Sackler School of Physics and Astronomy, Tel-Aviv University, Tel Aviv, 6997801, Israel.
Introduction. Non-Abelian phases of matter—ranging from fractional quantum Hall (FQH) states [1] to quantum spin liquids[2]—are at the forefront of strongly-correlated quantum many-body physics. Such phases are potential host of exotic quasiparticles exhibiting non-Abelian braiding statistics, which are proposed ingredients for topological quantum computation [3]. Half-integer or even-denominator FQH states [4] are a fertile playground for hosting such quasiparticles, thereby being a focal point of research over the past decades. The most appealing non-Abelian bulk phases [4] are the Pfaffian, its hole-conjugate partner—the anti-Pfaffian, and the particle-hole symmetric Pfaffian, discarding previously proposed Abelian platforms. Among the various experimental probes, thermal conductance offers profound insights into these phases [5, 6]. Whether dealing with bosons, simple integer quantum Hall states, or Abelian anyons [7, 8, 9, 10, 11, 12, 13, 14]— classified into simple particle-like and complex hole-like (with upstream neutral modes) FQH liquids — thermal conductance is consistently quantized to an integer value in the units of $\kappa_{0}T$, $\kappa_{0}=\frac{\pi^{2}k_{B}^{2}}{3h}$, where $k_{B}$ is the Boltzmann constant, $h$ is the Planck’s constant, and $T$ being the temperature. By contrast, a half-integer quantization of thermal conductance, $\frac{1}{2}\kappa_{0}T$, has been taken as a definite indicator of non-Abelian phases [4, 2], serving as compelling evidence for the presence of Majorana edge mode[5, 6]. Importantly, as has been demonstrated by earlier works[4], half integer values of the electrical conductance can emerge due to dynamics extraneous to the Abelian or non-Abelian nature of the underlying FQH phase.
A natural follow-up question is whether half-integer conductance values are exclusive to even-denominator FQH states. Interestingly, theoretical work [15] has proposed that fractional electrical conductance values can be engineered in graphene n-p or n-p-n junctions using integer QH states. Experimentally, both fractional and half-integer electrical conductance values have been observed in such systems [16, 17, 18, 19]. Moreover, half-integer electrical conductance has been realized in other platforms as well, such as quantum anomalous Hall insulator-superconductor junctions [20, 21, 22], and in quantum point contact (QPC) responses at $\nu=2/3$ in GaAs-based FQH systems [23, 24]. It is important to emphasize that these observations are not associated with non-Abelian bulk states but rather represent engineered half-integer electrical conductance in systems governed by Abelian physics. This raises a fundamental question: for such platforms that mimic fractional conductance plateaus, can one also expect a half-integer quantization of thermal conductance, $\frac{1}{2}\kappa_{0}T$, which is widely believed to be a hallmark of non-Abelian states. As of now, no theoretical or experimental studies have definitively addressed this question.
In this work, we demonstrate that half-integer thermal conductance can be realized without the presence of a Majorana mode. We utilize n-p-n heterojunctions consist of symmetry broken $\nu^{\prime}=-1$ region sandwiched between two $\nu=2$ integer QH state in the zeroth Landau level (ZLL) of bilayer graphene (BLG) to artificially create a $\frac{1}{2}G_{0}$ electrical conductance state, where $G_{0}=e^{2}/h$, is the electrical conductance quanta. The high tunability of symmetry-broken flavors (spin, valley, and orbital)[25, 26, 27, 28] in the ZLL of BLG — through both magnetic and electric fields — enables full equilibration between the co-propagating electron-hole (e-h) modes for which the two-terminal electrical conductance becomes, $G_{\nu-\nu^{\prime}-\nu}=\frac{|\nu||\nu^{\prime}|}{|\nu|+2|\nu^{\prime}|}G_{0}=\frac{1}{2}G_{0}$[17] (see "Methods" and supplementary). To test the idea that, two-terminal thermal conductance of such QH n-p-n junction also leads to a value of $\frac{1}{2}\kappa_{0}T$ (which is proposed by our theory as discussed in method and Supplementary Information (SI) section 12), we employ a unique device geometry with three arms as shown schematically in Fig. 1a. This geometry enables the creation of a temperature difference ($\Delta T$) across the n-p-n junction while maintaining zero net charge current, despite heat transport via chaotic (e-h) mixing across the junction. Our data suggests, full equilibration of both charge and energy leading to thermal conductance of $\sim\frac{1}{2}\kappa_{0}T$, consistent with our theoretical calculations. This challenges our current understanding as half-quantization of heat conductance is a unique signature of a Majorana mode, which makes this work is of fundamental importance. Our results also confirm that the Wiedemann-Franz law remains valid for engineered fractional values of thermal conductance, in contrast to the behavior observed in anyonic heat flow [11, 5, 13, 14]. We reproduce this result in another device consists of a p-n-p junction with $\nu=-2$ and $\nu^{\prime}=1$, reinforcing the reliability of our findings.
Figure 1: Device schematic, cross-section and optical image. (a) Device schematic and measurement set-up (see SI section 1 for more details). The device consists of three BLG arms connected by a central floating contact (FC), of which two identical arms on the right hand side are kept at $\nu=2$, controlled by global graphite back-gate and the other one having a local top-gate (green shaded region) in the middle with $\nu=-1$ forming an n-p-n segment with two interfaces of co-propagating electron-hole chiral-edges. The direction of the applied out of plane magnetic field dictates the chirality of the electron and hole edges, shown by the arrows associated with each chirals. The full charge and heat-equilibration along the length of the two interfaces is illustrated by the white glowing spheres. Each of the arms has atleast one cold ground contact labelled as CG. To measure the thermal conductance, noiseless dc currents $I_{S}$ and $-I_{S}$ are injected simultaneously at $S_{1}$ and $S_{2}$ respectively, which increases the electronic temperature of FC to $T_{M}$, while keeping it’s chemical potential (or voltage) to zero, and to determine the $T_{M}$, Johnson-Nyquist noise is measured both at $D_{1}$ and $D_{2}$, with the help of resonant LC network. The color of each chiral signifies it’s temperature. For the chirals emerging from FC are the hottest, shown by orange. Then after equilibration it reduces it’s temperature and hence shown by pink. The coldest ones are shown by blue, thermalized to temperature $T_{0}$. The hot-spot positions are marked by the fire signs which indicate the changing points for the chemical potential (or DC voltage) along the chirals. (b) Device cross-section showing the BLG is encapsulated by the top and bottom hBN. The global graphite back-gate and metallic local top-gate is connected external voltage source $V_{BG}$ and $V_{TG}$ respectively, which controls the respective filling factors. (c) Optical image of the device with the individual contacts marked same as in (a). The size of the scale bar is $5\mu m$.
Device and working principle. We use high-quality hBN encapsulated BLG heterostructures with graphite as the global back gate (BG) (see "Methods" for device fabrication details). As shown in Fig. 1a, the schematic of the device consists of three arms with a central floating contact (FC). To the right of the FC, there are two identical BLG channels with densities controlled by global BG. In contrast, in the left channel, a local top gate (TG) is placed over a specific region to create a n-p-n or p-n-p heterojunction. In this configuration, co-propagating electron and hole QH edge modes move along the interface of the n-p (and p-n) region, and their respective filling factors, $\nu$ and $\nu^{\prime}$ are tuned by $V_{BG}$ and $V_{TG}$. Here, our chosen platform BLG plays a crucial role in this experiment — not only it allows tuning between electron or hole carriers by adjusting the gate voltages, but it also facilitates equilibration between the e-h modes due to the tunability of symmetry-broken flavors (spin, valley, and orbital) of ZLL of BLG with magnetic and displacement field[28]. In the SI section 5, we show how the flavors of symmetry-broken edge modes of ZLL of BLG at the n-p (and p-n) interface depend on the displacement fields of the global (controlled by only BG) and local part (controlled by both BG and TG). If the flavors are same at the interface, one would expect full equilibration of chemical potential and energy, and the two-terminal electrical conductance across the junction is expected to be $G_{\nu-\nu^{\prime}-\nu}=\frac{|\nu||\nu^{\prime}|}{|\nu|+2|\nu^{\prime}|}G_{0}$[17] (see "Methods" for the theoretical calculation). For $\nu=2$ and $\nu^{\prime}=-1$, the $G_{\nu-\nu^{\prime}-\nu}$ expected to be $\frac{1}{2}G_{0}$, thus one can engineer half-integer electrical conductance. To measure thermal conductance across the n-p-n junction, our three-arm device (Fig. 1a) has the following advantages in contrast to our earlier works on two-arm devices[12, 13, 14, 29]. In order to create the temperature difference across the n-p-n junction, simultaneously, an electron and a hole current of equal magnitude are injected to the FC (from the identical two arms placed to the right side of the FC in Fig. 1a) to maintain its effective chemical potential at zero. This ensures no voltage drop across the n-p-n junction while having hotspots at the FC to increase its temperature as shown schematically in Fig. 1a. This is crucial because electrically biased graphene n-p-n juction would otherwise generate shot noise due to the current partitioning of co-propagating e-h modes at the n-p (and p-n) interface[30, 31], overshadowing the Johnson-Nyquist signal carrying the $\Delta T$ information. This device geometry enables us to overcome this experimental hurdle by eliminating shot noise as there is no voltage difference across the junction and therefore allows us to successfully measure the $\Delta T$ even though heat is carried through the chaotic e-h mixing across the n-p-n junction. For full thermal equilibration, the thermal conductance of such n-p-n junction will be, $K_{\nu-\nu^{\prime}-\nu}=\frac{|\nu||\nu^{\prime}|}{|\nu|+2|\nu^{\prime}|}k_{0}T$ (see "Methods" for the theoretical calculation) and for $\nu=2$ and $\nu^{\prime}=-1$, the $K_{\nu-\nu^{\prime}-\nu}$ expected to be $\frac{1}{2}\kappa_{0}T$ — a half-integer thermal conductance.
Figure 2: Device charaterzation. (a) Sketch of the measurement setup used to probe device response, using standard lock-in techniques. A current $I_{S}\sim 5nA$ is injected to either $S_{1}$ or $S_{2}$ while ensuring that at least one ground is present in each arm. The magnetic field direction is chosen such that, chirality of electronic carriers remain clock-wise. Black dashed line represents $I_{S}$ reaching the FC, where it splits into three parts. Two equal portions, labeled $I_{R}$ (blue dashed line) propagate along the right side of FC and terminate into respective CG, while the remaining one $I_{T}$ (red dashed line), travels along the left of FC, which is given by $I_{T}=V_{T}|\nu_{BG}|G_{0}$, where $V_{T}$ is the measured voltage drop at $D_{2}^{\prime}$ (see SI section 4 for more details). (b) 2D colormap of the measured $I_{T}$ normalized to $I_{S}$ as a function of $\nu_{BG}$ and $\nu_{TG}$. (c) Line cut along the black dashed-line in (b), showing the variation of $I_{T}/I_{S}$ with $\nu_{TG}$ while the bulk filling factor remains at $2$. Highlighted region around $\nu_{TG}=-1$ showing $I_{T}/I_{S}=0.11$ corresponding to $G_{T}=0.5G_{0}$. The vertical line divides at $\nu_{TG}=0$, separates the uni-polar and bi-polar region. The values of $I_{T}/I_{S}$ at other marked plateau regions, e.g., $\nu_{TG}=2$ and $1$, follows $G_{T}=min(|\nu_{BG}|,|\nu_{TG}|)G_{0}$.
Results:
Electrical conductance. To measure the electrical conductance across the n-p-n junction of our device geometry in Fig. 1a, we use the standard lock-in-based measurements with ac excitation frequency of $\sim 13Hz$. All the measurements are done in a cryo-free dilution fridge at base temperature, $T\sim 20mK$ (see "Methods"). After confirming the presence of well-developed symmetry-broken QH states at each arm (see SI section 3) of our device, we inject a constant current $I_{S}\sim 5nA$ at $S_{1}$ (or $S_{2}$) and measure the transmitted current through the n-p-n region (the top-gated part of the left arm of the device) in Fig. 1a. Upon reaching the FC, the injected current, $I_{S}$ divides into three parts, following to the conductance of each arm. These currents flow along the device boundaries and terminate at the respective grounds as shown schematically in Fig. 2a, which can also be understood in the simplified conductor model as shown in SI section 4. The conductance of the two identical arms on the right of the FC are same, and is given by $G_{R}=|\nu|G_{0}$, with corresponding current $I_{R}$. Whereas, for the left arm, the current, $I_{T}$ depends on the trans-conductance of the n-p-n junction, which follows $G_{T}=min(|\nu|,|\nu^{\prime}|)G_{0}$ in the uni-polar regime, and for the bi-polar regime, it is given by $G_{T}=\frac{|\nu||\nu^{\prime}|}{|\nu|+2|\nu^{\prime}|}G_{0}$ for the case of full charge equilibration of the co-propagating edges at the n-p (and p-n) interface. The current conservation follows as $I_{S}=2I_{R}+I_{T}$ and the ratio of $\frac{I_{T}}{I_{S}}=\frac{G_{T}}{2G_{R}+G_{T}}$. Since, $G_{R}$ is known, hence, by measuring the transmitted current $I_{T}$ through the n-p-n junction, one can infer the $G_{T}=G_{\nu-\nu^{\prime}-\nu}$. The details about $I_{T}$ measurement are discussed in the SI section 4.
The measured $I_{T}/I_{S}$ as a function of $\nu_{BG}$ ($\nu$) and $\nu_{TG}$ ($\nu^{\prime}$) is shown as a 2D colormap in Fig. 2b for 6T magnetic field. The vertical dark blue strip around $\nu_{TG}=0$, where $I_{T}=0$, indicates the top-gated region becomes insulating, making the boundary between the uni-polar and bi-polar region. Around $\nu_{BG}=0$ we also observe $I_{T}=0$ as the entire bulk becomes insulating. The right-half of the plot, representing the uni-polar region, is divided into rectangular blocks of different colors centralized around the integer values of $\nu_{BG}$ and $\nu_{TG}$ signifying the robust plateaus of $I_{T}$ at different filling configurations as expected. For instance, for $\nu_{BG}=\nu_{TG}$, the injected current $I_{S}$ is expected to divide equally among the three arms leading to $\frac{I_{T}}{I_{S}}=0.33$. This is indeed observed in the measured data for $(\nu_{BG},\nu_{TG})=(1,1)$, $(2,2)$, and $(4,4)$, and also close to the expected value for $(3,3)$ as well. This also satisfies an essential requirement for thermal conductance measurements – equipartition of current at FC, and will be discussed later in details (see the simplified conductor model illustrated in SI section 4). However, in the bi-polar region (left-half portion), the $\frac{I_{T}}{I_{S}}$ is not simple for different combinations of ($\nu_{BG},-\nu_{TG}$) since its value depends on the degree of equilibration of the co-propagating edges at the n-p (and p-n) interface. It can be seen from Fig. 2b (white dotted boxes) that for (2,-1) and (1,-2) the measured $I_{T}$ suggests robust plateau with full equilibration ($G_{\nu-\nu^{\prime}-\nu}=\frac{|\nu||\nu^{\prime}|}{|\nu|+2|\nu^{\prime}|}G_{0}$) in contrast to (1, -1) and (2,-2) with partial equilibration (measured value $<\frac{|\nu||\nu^{\prime}|}{|\nu|+2|\nu^{\prime}|}G_{0}$). The full and partial equilibration can be understood in terms of the spin state of the co-propagating edges at the interface, as discussed in the SI section 5 based on prior studies[32, 28], and also present the data with various magnetic fields revealing a clear transition of (2,-1) from partial to full equilibration. In Fig. 2c, we show the trace of $\frac{I_{T}}{I_{S}}$ as a function of $\nu_{TG}$ while $\nu_{BG}=2$. It can be seen for (2,-1), $\frac{I_{T}}{I_{S}}\sim 0.11$ with robust plateau indicating $G_{\nu-\nu^{\prime}-\nu}=\frac{|\nu||\nu^{\prime}|}{|\nu|+2|\nu^{\prime}|}G_{0}=0.5G_{0}$. This is repeated in another device, as shown in SI section 9. Though, the equilibration for many combinations of filling factors can be understood in terms of the spin state of the co-propagating edges, however, anomalously suppressed equilibration observed for (2,-2) requires further theoretical understanding beyond the spin-selective equilibration model (see SI section 5 for the discussion).
Figure 3: Thermal-Hall conductance in the uni-polar region. (a)-(c) Excess thermal noise is measured at $D_{1}$ as a function of injected dc current $I_{S}$ for different configuration of $(\nu_{BG},\nu_{TG})=(2,0)$ (a), $(2,1)$ (b) and $(2,2)$ (c). (d) The increased temperatures $T_{M}$ of floating contact, are extracted from panel (a)-(c), are plotted (solid circles) against the dissipated power $J_{Q}$($=P$), due to Joule heating in FC. (e) $J_{Q}/0.5\kappa_{0}$ is plotted (solid circles) as a function of $(T_{M}^{2}-T_{0}^{2})$ for $(2,2)$, $(2,1)$ and $(2,0)$. The solid lines are the linear fittings to extract the thermal conductance, parameterized as $N\kappa_{0}$ (Eq. 1), with $N$ as the of outgoing channels. Slope of this fits give $N$ as $6.04$, $5.01$ and $4.03$, respectively as expected. A simple diagram of the heat-flow depicted in the right. From this we can deduce $N=2|\nu_{BG}|+min(|\nu_{BG}|,|\nu_{TG}|)$.
Thermal conductance. To measure the thermal conductance, we simultaneously inject dc currents, $I_{S}$ and $-I_{S}$ into FC from sources $S_{1}$ and $S_{2}$, respectively, as shown in Fig. 1a to maintain the effective chemical potential of FC at zero. In this configuration, the power dissipation at FC is given by $P=\frac{I_{S}^{2}}{\nu_{BG}G_{0}}$[12, 13, 14, 29] (See "Methods" for details). This results in an increase in the electronic temperature of the FC and, its steady-state temperature, $T_{M}$ is determined by the following heat balance relation[10]:
$P=J_{Q}=J_{Q}^{e}(T_{M},T_{0})+J_{Q}^{e-ph}(T_{M},T_{0})=0.5N\kappa_{0}(T_{M}^{2}-T_{0}^{2})+J_{Q}^{e-ph}$ (1)
Here, $J_{Q}^{e}(T_{M},T_{0})$ represents the electronic contribution of the heat current via $N$ outgoing chiral channels from the FC, $J_{Q}^{e-ph}$ accounts for the heat loss due to electron-phonon cooling, and $T_{0}$ is the electron temperature of the cold reservoirs. To obtain $T_{M}$, we measure the excess thermal noise at detectors $D_{1}$ (or $\bar{D_{1}}$) and $D_{2}$, positioned along the outgoing edge channel and using the Nyquist-Johnson relation[10, 11], $S=2G^{*}k_{B}(T_{M}-T_{0})$ for uni-polar case (i.e., when $\nu_{BG}$ and $\nu_{TG}$ have the same sign). Here, $\frac{1}{G^{*}}=\frac{1}{G_{amp}}+\frac{1}{\sum_{i=1,i\neq amp}^{3}{G_{i}}}$, $G_{amp}$ is the conductance of the arm with the detector, and $G_{i}$ represents the conductance of other arms. For the right arms in Fig. 1a, the conductance, $G_{i}=\nu_{BG}G_{0}$, while for the left arm, $G_{i}=min(\nu_{BG},\nu_{TG})G_{0}$.
Before we present the thermal conductance data for the bi-polar regime, we first benchmark the thermal conductance values for the uni-polar region. As shown in Fig. 2c, we see robust plateaus around $\nu_{TG}=0$, $1$ and $2$, while the bulk filling is kept at $\nu_{BG}=2$. We measure the excess noise, $S_{I}$, at the detector $D_{1}$ for (2,0), (2,1) and (2,2), and shown in Fig. 3a-c, as a function of injected dc current $I_{S}$ (and $-I_{S}$). The extracted $T_{M}$ from the $S_{I}$ for different $(\nu_{BG},\nu_{TG})$ are shown in Fig. 3d, against the dissipated power, $J_{Q}$. The electronic temperature $T_{0}$ is $\approx 20mK$ (its determination is shown in SI section 2). In Fig. 3e, the $J_{Q}/0.5\kappa_{0}$ is plotted with $(T_{M}^{2}-T_{0}^{2})$. The linear fitting of each plot will give the value of the thermal conductance, and the extracted values are $6.04\kappa_{0}$, $5.01\kappa_{0}$ and $4.03\kappa_{0}$, respectively, for $(\nu_{BG},\nu_{TG})=(2,2)$, $(2,1)$ and $(2,0)$. The linearity of each plot suggests a negligible e-ph contribution (second term in Eq. 1) within the displayed range of $T_{M}\approx 50mK$. Our results are consistent with the simplified heat-flow diagram shown in the right of Fig. 3e: when $\nu_{BG}=2$, two identical right arms of the FC always hosts total $4$ number of channels, while the number of channels, on the left arm of the FC, are varied from $2$ to zero by tuning $\nu_{TG}$, thus, in general, one can express the total number of out-going channels ($N$) from the FC as $N=2|\nu_{BG}|+min(|\nu_{BG}|,|\nu_{TG}|)$.
Figure 4: Half-integer thermal conductance. (a) Excess thermal noise measured at the half-conductance plateau $(\nu_{BG},\nu_{TG})=(2,-1)$, at $D_{1}$ (red circles, $S_{D1}$) and at $D_{2}$ (black Circles, $S_{D2}$) as a function of $I_{S}$. $S_{Linear}=S_{D1}-\frac{1}{4}S_{D2}$ is shown by the blue circles. (b) Inset: Increased temperature of the floating contact $T_{M}$ extracted from $S_{Linear}$ in panel (a) is plotted against the dissipated power $J_{Q}$. $\Delta J_{Q}/0.5\kappa_{0}$ is plotted (solid circles) as a function of $T_{M}^{2}-T_{0}^{2}$. The slope of the linear fitting (solid line) gives the thermal conductance value which is expected to be $4.5\kappa_{0}$, which is depicted by the heat-flow diagram having $4$-channel of heat conductance of $\kappa_{0}$ and one effective channel with net thermal conductance of $0.5\kappa_{0}$. (c) $\Delta J_{Q}/0.5\kappa_{0}$ is plotted against $T_{M}^{2}$ for $\Delta N=2$ (between $[2,2]$ $[2,0]$), $\Delta N=1.5$ (between $[2,2]$ $[2,-1]$), $\Delta N=1$ (between $[2,2]$ $[2,1]$) and $\Delta N=0.5$ (between $[2,-1]$ $[2,0]$), where $\Delta J_{Q}=J_{Q}(N_{i},T_{M})-J_{Q}(N_{j},T_{M})$. The solid lines are linear fittings to extract the the thermal conductance associated with $\Delta N$ $1D$ channels. Inset shows the slopes extracted from each solid lines vs $\Delta N$. The error bar represents the uncertainty of the fitting. The dashed line shows the theoretical $1\kappa_{0}$ line.
Half-integer thermal conductance. In this section, we present the thermal conductance for the bi-polar regime of $(\nu_{BG},\nu_{TG})=(2,-1)$. However, generically in the bi-polar regime, the noise [33, 34, 35, 36] to temperature conversion is not simple and requires solving higher-order equations (see the "Methods" and the SI section 12 for details). A simplified approach is to measure the excess noise simultaneously both at $D_{1}$(or $\bar{D_{1}}$) and $D_{2}$, leading to a linear relation (see SI section 12 for details):
$S_{Linear}=S_{D1}-\frac{1}{4}S_{D2}=|\nu_{BG}|G_{0}k_{B}(T_{M}-T_{0}).$ (2)
For the (2,-1) plateau, the excess thermal noise, $S_{D1}$ (red circles) and $S_{D2}$ (black circles) measured at $D_{1}$ and $D_{2}$, respectively, are shown in Fig. 4a as a function of $I_{S}$ (and $-I_{S}$). The $S_{Linear}$ is shown by the blue circles in Fig. 4a, which is further converted to $T_{M}$ using Eq. 2 and shown as a function of $J_{Q}$ in the lower inset of Fig. 4b. The $J_{Q}/0.5\kappa_{0}$ is plotted (blue circles) as a function of $T_{M}^{2}-T_{0}^{2}$ in Fig. 4b and the linear fitting gives the value of the thermal conductance $\sim 4.55\kappa_{0}$, which is depicted by the heat-flow diagram having $4$ out-going channels via the two identical right arms (in Fig. 1a), and an effective channel with net thermal conductance of $\sim 0.5\kappa_{0}$ through the left arm consists of the n-p-n junction ($(\nu_{BG},\nu_{TG})=(2,-1)$). This suggests the thermal conductance of the n-p-n junction is $\sim 0.5\kappa_{0}$. To further confirm it, we plot the $\Delta J_{Q}/0.5\kappa_{0}$ with $T_{M}^{2}$ for different configurations of $\Delta\nu$: $\Delta$([2,2] - [2,0]), $\Delta$([2,2] - [2,-1]), $\Delta$([2,2] - [2,1]) and $\Delta$([2,-1] - [2,0]) in Fig. 4c, and the corresponding linear fittings give the thermal conductance values of $\sim 2.03\kappa_{0}$, $1.45\kappa_{0}$, $0.95\kappa_{0}$, $0.54\kappa_{0}$, respectively. As expected for $\Delta$([2,2] - [2,0]) and $\Delta$([2,2] - [2,1]), the thermal conductance values are consistent with the effective number of channels $\Delta N=2$ and $\Delta N=1$, respectively. However, $\Delta$([2,2] - [2,-1]) with $\sim 1.45\kappa_{0}$ and $\Delta$([2,-1] - [2,0]) with $\sim 0.54\kappa_{0}$ reconfirms the half-integer thermal conductance of the n-p-n junction with the effective number of channels $\Delta N=0.5$. This is further reemphasized in Fig. 4c inset via plotting the extracted thermal conductance values with $\Delta N$, and we see a linear increment with the slope of $1\kappa_{0}$. Similar observations are presented for Device 2 in Supplementary Information Section 11.
Discussion and outlook. The observation reported here of a plateau of half-integer thermal (as well as electrical) conductance relies not on underlying topology characterizing non-Abelian phases, but rather on the robustness of fully-equilibrated edge modes to local perturbations. Specifically, we address a "bi-polar" boundary, (2,-1), separating two integer quantum Hall phases. The feasibility of inter-mode full equilibration in our study depends on the flavors of symmetry-broken edge modes as well as on the presence of disorder[37, 38, 39, 15] at the bipolar junction. Our device, having a boundary of length $\sim 10\mu m$, an order of magnitude larger than earlier works[37, 40, 41], allows full equilibration of edge modes. The resulting edge transport is expected to be incoherent for our extended transport path, consistent with earlier findings[31], where the measured shot noise Fano distinguishes the coherent versus incoherent process at the bipolar regime.
Our work puts on the table the issue of robustness of quantum transport platforms to external perturbations. One may compare topology-based resilience to resilience underlined by strong equilibration dynamics. Such a study can use case studies of other equilibration-engineered fractional values of thermal and electrical conductance, as well as also serving a test-bed for the validity of the Wiedemann-Franz law. This might be particularly interesting when implemented to platforms which (in the absence of equilibration) exhibit charge fractionalizaton giving rise, e.g., to charged and neutral eigenmodes [42, 43, 44, 45]. Generalizations to fractional quantum Hall boundaries (e.g., with charge and neutral modes [46, 44] or charge and spin modes [47]) should follow. This work also opens up an intriguing direction of incorporating equilibration dynamics (e.g. [48]) into the physics of dilute colliding beams[49, 50, 51, 52], involving both fermions and anyons.
Note added– While completing this draft, we came to know of a recent manuscript by Karmakar et al.: arXiv: 2505.08746 (2025).
Methods
Device fabrication. Our devices consists of Van-der Waals heterostructure of hBN/BLG/hBN/Graphite on a p-doped Si substrate with $285nm$ of $SiO_{2}$ and was fabricated using conventional dry-transfer technique[53]. All the Ohmics including the floating contact were defined using electron-beam lithography (EBL) followed by reactive-ion etching with $CHF_{3}$ and $O_{2}$. After that Cr/Pd/Au $(5/12/60nm)$ were deposited using a thermal evaporator kept at a pressure of $1-2\times 10^{-7}$ mbar. The local top-gate was then fabricated with a subsequent EBL and metallization. Finally, the device geometry was patterned with EBL and etched using $CHF_{3}$ and $O_{2}$.
Measurements. All the measurements were carried out in a cryo-free dilution refrigerator with a base temperature of $\sim 20mK$. The transport measurements were performed using standard lock-in techniques. The filling conversions from the applied gate voltages $V_{BG}$ and $V_{TG}$ at a particular magnetic field $B$, are given as follows:
$\nu_{BG}=\frac{C_{BG}\phi_{0}}{B}(V_{BG}-V_{BG0})\quad\text{and}\quad\nu_{TG}=\frac{C_{BG}\phi_{0}}{B}(V_{BG}-V_{BG0})+\frac{C_{TG}\phi_{0}}{B}(V_{TG}-V_{TG0})$
Here, $\phi_{0}=\frac{h}{e}$ is the magnetic flux quanta and $V_{BG0}$ and $V_{TG0}$ are the positions the charge neutrality point (CNP) of the back-gated and top-gated regions, respectively.
For noise measurements, we use a resonant $LC$-tank circuit tuned around $\sim 750kHz$. The signal was amplified using a homemade cryo-amplifier based on HEMT, operating at $4K$. This is followed by further amplification at room temperature, and finally recorded using a spectrum analyzer (see detailed schematic in the SI section 1). At $I_{S}=0$, the amplifier measures the equilibrium voltage fluctuation given by,
$S_{V}(0)=G^{2}(4k_{B}TR+v_{n}^{2}+i_{n}^{2}R^{2})BW$ (3)
Here, $G$ is the total gain of the amplifier chain, $BW$ is the measurement bandwidth, and $v_{n}$ and $i_{n}$ are the voltage and current noise of the amplifier, respectively. The first term in the equation represents the thermal noise due to the finite electronic temperature of the system. At finite bias, Joule heating in the FC elevates its temperature, adding excess thermal noise on top of $S_{V}(0)$. This excess noise is calculated by subtracting the zero-bias noise, $\delta S_{V}=S_{V}(I_{S})-S_{V}(0)$. Finally, the excess current noise is obtained by, $S=\frac{\delta S_{V}}{R^{2}}$, where $R$ is the resistance seen by amplifier, given by $R=\frac{h}{\nu_{BG}e^{2}}$.
Theoretical model: We refer to Fig. 1a and assume full charge and thermal equilibration among different edges allowing us to define local voltage and temperature at each junction of the device. Henceforth, we theoretically calculate the following quantities to analyze the transport and excess noise in the device in the n-p-n region (see supplementary for details). Notably, the calculations are also valid for the p-n-p region and any effect of edge reconstruction is washed out because of full equilibration.
(1) Electrical conductance– We assume full charge equilibration in each segment of the device (c.f. Fig. 1a) leading to $L\gg l^{\text{ch}}_{\text{eq}}$, where $L$ and $l^{\text{ch}}_{\text{eq}}$ are the geometric length and charge equilibration length respectively. To calculate the electrical conductance plateau, we ground the contact $S_{2}$ and bias the contact $S_{1}$ by a dc voltage $V_{0}$ corresponding to a source current $I_{S}=|\nu_{\text{BG}}|V_{0}\frac{e^{2}}{h}$. Electrical current conservation at each junction of the device results in the following.
$I_{D_{2}}=\frac{|\nu_{\text{BG}}||\nu_{\text{TG}}|}{2|\nu_{\text{BG}}|+5|\nu_{\text{TG}}|}V_{0}\frac{e^{2}}{h},I_{D_{1}}=I_{\bar{D}_{1}}=\frac{|\nu_{\text{BG}}|(|\nu_{\text{BG}}|+2|\nu_{\text{TG}}|)}{2|\nu_{\text{BG}}|+5|\nu_{\text{TG}}|}V_{0}\frac{e^{2}}{h},$ (4)
where $I_{D_{2}},I_{D_{1}},I_{\bar{D}_{1}}$ are the electrical currents in drains $D_{2},D_{1},\bar{D}_{1}$ respectively, satisfying $I_{S}=I_{D_{2}}+I_{D_{1}}+I_{\bar{D}_{1}}$. The potential $V_{M}$ of the FC is
$V_{\text{M}}=\frac{|\nu_{\text{BG}}|+2|\nu_{\text{TG}}|}{2|\nu_{\text{BG}}|+5|\nu_{\text{TG}}|}V_{0}$ (5)
and thereby the two-terminal electrical conductance $G_{\text{T}}$ corresponding to the n-p-n region is
$G_{\text{T}}=\frac{I_{D_{2}}}{V_{\text{M}}}=\frac{|\nu_{\text{BG}}||\nu_{\text{TG}}|}{|\nu_{\text{BG}}|+2|\nu_{\text{TG}}|}\frac{e^{2}}{h}.$ (6)
Now for $|\nu_{\text{BG}}|=2,|\nu_{\text{TG}}|=1$, we have
$\frac{I_{D_{2}}}{I_{S}}=\frac{1}{9},\ G_{\text{T}}=\frac{1}{2}\frac{e^{2}}{h}$ (7)
consistent with the experimental data in Fig. 2c.
(2) Thermal conductance– To calculate the thermal conductance, we bias both the contacts $S_{1}$ and $S_{2}$ corresponding to a source current $I_{S}$ in $S_{1}$ and $-I_{S}$ in $S_{2}$ (c.f. Fig. 1a). This makes the effective electrochemical potential of the FC to be zero. Now, the heat balance equation dictates that the total power ($P$) dissipated at the floating contact is equal to the total heat current ($J_{Q}$) flowing out of it and therefore $P=J_{Q}$. Voltage drops at the FC lead to $P=|\nu_{\text{BG}}|e^{2}V_{0}^{2}/h$. We denote the number of co-propagating edge modes for $|\nu_{\text{BG}}|$ is $|n|$ and that of $|\nu_{\text{TG}}|$ is $|m|$. In the full thermal equilibration regime, leading to $L\gg l^{\text{th}}_{\text{eq}}$ where $l^{\text{th}}_{\text{eq}}$ is the thermal equilibration length, we have
$J_{Q}=\frac{\kappa_{0}}{2}\kappa_{\text{th}}(T_{M}^{2}-T_{0}^{2}),\ \kappa_{\text{th}}=\Bigg{[}2|n|+\frac{|m||n|}{2|m|+|n|}\Bigg{]},$ (8)
where $T_{M}$ is the temperature of the FC and $T_{0}$ is the temperature of all other contacts. The term $\frac{|m||n|}{2|m|+|n|}$ arises solely due to the n-p-n region. Notably, for the particular case considered here, we have $\kappa_{\text{th}}=4.5$ for $|n|=2,|m|=1$ consistent with the experimental data in Fig. 4b.
(3) Johnson-Nyquist noise– To compute $T_{M}$ experimentally, a known prescription is to use the excess noise. Therefore, we calculate the excess noise $S_{D_{1}}(=S_{\bar{D}_{1}})$ and $S_{D_{2}}$ at drains $D_{1}$ (or $\bar{D}_{1}$) and $D_{2}$, respectively, by using the following procedure. At each junction of the device, we write the conservation equations of electrical current fluctuation and each of which has two contributions—a fluctuation from local voltage and a thermal fluctuation. The latter is related to the local temperature via Johnson-Nyquist noise. Now, we self-consistently solve these equations to calculate the excess noise at each drain. Our calculations show that $S_{D_{1}}(=S_{\bar{D}_{1}})$ and $S_{D_{2}}$ depend on $T_{0}$ and $T_{M}$ in complicated ways as
$\begin{split}&S_{D_{1}}(=S_{\bar{D}_{1}})\sim f_{1}\Big{(}T_{0},T_{M},\sqrt{T^{2}_{0}+T^{2}_{M}}\Big{)}\frac{e^{2}k_{B}}{h},\\ &S_{D_{2}}\sim f_{2}\Big{(}T_{0},T_{M},\sqrt{T^{2}_{0}+T^{2}_{M}}\Big{)}\frac{e^{2}k_{B}}{h},\end{split}$ (9)
where $f_{1}$ and $f_{2}$ depict distinct functional dependence on $T_{0},T_{M}$ (the exact form is shown in the supplementary). Extracting $T_{M}$ from those requires to solve higher-order equations. Remarkably, we find that the a suitable combination ($S_{\text{Linear}}$) of $S_{D_{1}}$ and $S_{D_{2}}$ leads to the following expression where $S_{\text{Linear}}$ is linearly dependent on $T_{M}$ and therefore
$S_{\text{Linear}}=S_{D_{1}}-\frac{1}{4}S_{D_{2}}=|\nu_{\text{BG}}|\frac{e^{2}}{h}k_{B}(T_{M}-T_{0}),$ (10)
as shown in the experimental data in Fig. 4a.
Data availability
The data presented in the manuscript are available from the corresponding author upon request.
Acknowledgements
U. R. and S. M. thanks Arup Kumar Paul for useful discussions. We thank the International Centre for Theoretical Sciences (ICTS) for participating in the program - Condensed Matter meets Quantum Information (code: ICTS/COMQUI2023/9), where the collaboration was initiated. S.M. was supported by the Weizmann Institute of Science, Israel Deans fellowship through Feinberg Graduate School. Y. G. was supported by the InfoSys Chair, IISc, Bangalore. S.M. and Y. G. were also supported by the Minerva Foundation and grant no 2022391 from the United States–Israel Binational Science Foundation (BSF), Jerusalem, Israel. Ankur Das was supported by IISER, Tirupati Startup grant and ANRF/ECRG/2024/001172/PMS. M.G. has been supported by the Israel Science Foundation (ISF) and the Directorate for Defense Research and Development (DDR&D) through Grant No. 3427/21, the ISF Grant No. 1113/23, and the US-Israel Binational Science Foundation (BSF) through Grant No. 2020072. A.D. thanks the Department of Science and Technology (DST) and Science and Engineering Research Board (SERB), India, for financial support (SP/SERB-22-0387), (DST/NM/TUE/QM-5/2019), and also acknowledges funding through the Intensification of Research in High Priority Areas programme of the Science and Engineering Research Board (Grant No. IPA/2020/000034). A.D. also thanks CEFIPRA project SP/IFCP-22-0005. Growing the hBN crystals received support from the Japan Society for the Promotion of Science (KAKENHI grant nos. 19H05790, 20H00354 and 21H05233) to K.W. and T.T. The authors gratefully acknowledge the use of Blender (https://www.blender.org/) and Inkscape (https://inkscape.org/) — both free and open- source software —- for the creation and refinement of figures presented in this work.
Competing interests
The authors declare no competing interests.
See pages 1-18 of SI.pdf